OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 3x − 2y = 6. x-intercept: (0, −3), y-intercept: (0, 2) x-intercept: (−3, 0), y-intercept: (2, 0) x-intercept: (0, 2), y-intercept: (0, −3) x-intercept: (2, 0), y-intercept: (0, −3)

OpenStudy (ash2326):

why. math see my answer in you previous question and try to do this one. If you get stuck, do let me know. Thanks:)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i dont get it

OpenStudy (ash2326):

did you understand the standard equation ?

OpenStudy (ash2326):

we are given the equation 3x-2y=6 now divide this by 6 \[\frac{x}{2}+\frac{-y}{3}=6\] now let's take the minus sign to the denominator \[\frac{x}{2}+\frac{y}{-3}=6\] so the x intercept is (2, 0) and y intercept is (0, -3) option D

OpenStudy (anonymous):

Thanks for explaining i understand the concept but i still dont actually understand it if the question were to be rephrase i wouldnt be able to answer it or even thoe i know exactly what to do.